mainmenu

DFR Telecoms Diary

2021

6 January : Posted by Rick

Welcome to 2021

Only Rick in today. James in shop today so checked Norchard exchange, passed test calls and tested junctions. Unfortunately the tester was not working. One of the tester batteries had gone open cct and the rest were low voltage so all the batteries were taken away for renewal. Not sure when they will be replaced in tester as we are now in lockdown.

8 January : Posted by Rick

A sad day today as Ray Willey has died of the Corona virus - he was a valued member of the team and our clock and PA specialist. He did like a good whinge but was always cheerful and ready with a humorous story from his past. He will be greatly missed.

13 January : Posted by Rick

Rick in today to liaise with Chris Bull and look at the tower lights at the coal stack and the steam train siding. We used a ladder and Chris used his tree climbing safety gear to reach the top of the siding tower after we had disconnected the electric. He discovered that one light was disconnected (probably by Alex Renny when he fitted the one existing led floodlight and thought that there was insufficient electric current to power three floodlights) but he could not access the light bulbs because the cover bolts were rusted on. He did manage to disconnect the wiring and undo the lighting fixing bracket and lower the whole unit. On further inspection on the ground it was decided that it was too badly corroded and should be renewed. We tidied up and reconnected the power and checked that the one floodlight was still working. Then decided that as it was raining we would leave the tall tower for another day.

13 January : Posted by Rick

Very bad weather at this time - cold or wet and windy - basically office tidied and Norchard exch tester batts prepared for refitting.

15 January : Posted by Rick

Zoom meeting this evening where the cable damage at Parkend was discussed and it was decided that there was no money available to purchase a new length of cable but if new cable was provided it would need to be laid in troughing - some troughing existed but was about 75 metres too short.

28 January : Posted by Rick

Alastair Clarke e mailed me to say they had some money and we can order the cable.

28 January : Posted by Rick

I spoke to Roger Phelps who said he would order the cable.

28 January : Posted by Rick

Roger Phelps rang to say the Norchard to Lydney Junction token did not work - could I test the batteries? A job I had not done before. I went to Lydney Junction signal box but was unable to check the batteries as there was no key for the battery cupboard. Then to Norchard but again no cupboard key available. Fortunately I met Steve Harris who lent me his key and I was able to test the batteries at both ends. All testing ok and passed back to Roger for further testing. I have now found a key on a bunch that John Bathgate returned to me so next time I will be prepared.

29 January : Posted by Rick

We had a zoom meeting in the evening as it had been some time since we had all been together. Rick Peter Martin Charles and Alan on screen and a lively discussion took place although zoom meetings are always a bit awkward and disjointed.

30 January : Posted by Rick

Sam rang me to say that the missing 20 pr/0.9 cable had been found. It seems it had been put at the back of the shop building and then a some wood had been stacked in front of it and now the wood had been used up so the cable had reappeared. We hope to make good use of it at Parkend.

1 February: Posted by Rick

Sam rang me to say the 50pr/ 0.9 cable had arrived at Norchard and had been put next to the other cable at the back of the shop building. Janet and I went to Norchard to check both cables and to cover them with tarpaulins and tie them down until we get around to using them. Someone is going to be very busy in the next few months.

4 February: Posted by Rick

Ray Willey’s funeral took place at Bournemouth today. Although we were unable to attend some of us watched as it was streamed over the internet from the graveside.

10 February: Posted by Rick

Rick and Janet in today and went to Parkend first to check and measure a possible site for a Panasonic switch to interface strowger / digital. While there the test call sender was run and 30 test calls no failures. We returned to Norchard and met Martin for tea/coffee (socially distanced of course) and a chat. Met Steve Harris and discussed Parkend signal box - hopefully completion October 21 or Easter 22. Spoke to Peter on the digital network from our office about Parkend..

17 February: Posted by Rick

Rick and Janet in today. We went to Whitecroft and spent the morning recovering 50 metres of 10pr/0.9 cable. Old cable along the fence from DP 3Z but tested ok. It should be useful somewhere as we have so many cable jobs coming up. The weather was good and it needed recovering before the brambles start growing. Now at Norchard for safety! The new batteries for Norchard exchange tester are now ready just need access..

18 February: Posted by Rick

Rick in today and arranged access to Norchard exchange to fit new batteries in the tester. Batteries fitted ok. Tested the Parkend junctions and all showing various amounts of low earth or low insulation. Need to check these further in order to be ready for the new Parkend cable..

22 February: Posted by Rick

Rick met Chris Bull at Norchard after lunch and they fitted the new floodlights to the lighting tower at the back of the workshop area ( Chris did all the climbing and fitting - Rick just passed him the spanners) and they all worked. A good bit of inter department cooperation!.

24 February: Posted by Rick

Rick and Janet in. Signal post telephone changed at the stop board opposite Norchard signal box and tested ok. We then walked the line to Middle Forge. When the Middle Forge project is completed it may be possible to pull the slack cable back to DP 3F and eliminate the tin can joint which would remove a point of possible cable damage. We collected some telephone pole furniture on our walk in the hope that we may be able to do a display in the museum. We also tested the recovered cable with the tester 301 - 66 metres to a clean dis so could be useful. Steve Harris dropped of an empty cable drum for the recovered cable..

1 March: Posted by Rick

Peter and Rick met Steve Harris at Parkend to discuss how many SPTs are required. We walked the line to check the plan to move cables and DPs. All good only need time and volunteers to complete..

3 March: Posted by Rick

Rick and Janet in today. One and a half hours spent preparing a weatherproof phone to replace the one at Norchard signal post 2. The existing coffin type changed to a modern type. Very difficult to fit to side of cabinet especially as the cabinet is at an angle of about 20 degrees. After fitting we discovered there was a line fault (very noisy). After a late lunch we checked the line but there were so many joins in the dropwire we decided to to tackle it on another day..

4 March: Posted by Rick

Rick and Janet in today to fault signal post 2. We eventually found a place where the dropwire had been crushed and the internal wires damaged. This was repaired but the dropwire is in such a poor state it really wants renewing. Just another ten minute job to add to the list..

10 March: Posted by Rick

Rick and Janet in today - hopefully the troughing will be delivered today. We made a temporary test point and tested 3 SPTs ready for use. Early home today..

17 March: Posted by Rick

Rick Janet and Peter in today. We all went to Lydney Junction as S and T have requested a new SPT at the south end somewhere. Guess they want it down at the point where the mainline enters. Back to Norchard and a bit more tidying in the office (might finish this job one day)..

18 March: Posted by Rick

Roger Phelps rang me at home at 14.00 hrs to ask why the broadband at Parkend was not working? I have no idea but I went to Parkend to check (it was time to take the dog for a walk anyway). It all looked ok and I could use the internet ok (another RWT) and the dog had a good run..

20 March: Posted by Rick

Rick and Peter met Steve Harris at Lydney junction to see what is required. A new SPT is needed where the main line joins DFR rails and the point is operated as the signal box is out of sight. We will draw up a works schedule to list stores required and costs and pass it to Steve.

21 March: Posted by Rick

Today the troughing at the Parkend straight was successfully laid by a team organised by Dan Winter and Alastair Clarke. This allows the cable to be laid which will replace the one damaged in June 2019 and allow communications to be restored to Whitecroft and Parkend.

24 March: Posted by Rick

Arranged to meet Peter, Dave Collins and Roger Phelps at Parkend to sort out the broadband problem. It works ok around the platform area but not in the signal box. It is hoped to bring the signal box into operation at the end of the year and Roger needs broadband access in order to have remote access to the new axle counter equipment. (ASAP of course) In order for this to be possible a new cat 5 cable will need to be provided from the data equipment in Parkend exch into the signal box along with the associated ancillary equipment (sockets cords etc). A plan of action was agreed - Peter will give Roger a list of requirements and costs who will arrange finance.

28 March: Posted by Rick

Today is the day the replacement cable is going to be laid at the Parkend straight which will allow the cable which was damaged in June 2019 to be replaced with a new 445 metre length of 50pr/0.9. The team was again organised by Dan Winter. They followed instructions and worked together so well that laying the cable in the troughing and replacing the lids (where available) was completed in just 1 hour. That is ignoring the time spent organising the train loading the cable assembling the team and travelling to site. All that is needed now is to join the new to the old and hope it still all works.

29 March: Posted by Rick

Rick and Janet stripped the outer and inner insulation of the new cable at DP1E and fanned out the wires ready for reconnecting.

30 March: Posted by Rick

Rick and Janet repeated yesterday’s work at DP1F.

31 March: Posted by Rick

Rick Peter and Charles in today at Parkend. Norchard to Parkend token cct changed over to new cable. Tested working ok by Dave Collins at Parkend and Roger Phelps at Norchard.

3 April: Posted by Rick

Rick and Peter at Parkend today tracing junctions (morning only).

4 April: Posted by Rick

Rick and Janet in tracing faulty junction between Tufts bridge and Middle forge. Possible damage by rats somewhere near Tufts. Requires further investigation.

7 April: Posted by Rick

Rick Peter and Charles in today at Norchard. Weather very cold today. Unable to visit Parkend as BBC filming in progress until Friday. Charles was sorting some of the electrical equipment wiring and fittings. Peter was collecting data equipment and fittings ready for Parkend. Rick was able to access and test ccts in Norchard exchange but in the end the cold won and we went home early.

10 April: Posted by Rick

Rick and Peter at Parkend to cut holes in the exchange floor ready for new cables. Unfortunately the film crew is still filming and did not want any extraneous noises to spoil it. Had to wait until lunch (13.00 - 14.00). Peter did some preparatory work and Rick went to Whitecroft and reconnected 308 in the station building. During lunch we managed to drill two holes in more or less the right places so now waiting for cable.

11 April: Posted by Rick

Rick and Janet went to Tufts bridge and used the tester 301 to look at the faulty junction. We tested from DP 3X towards Norchard. It showed a low a to b at 9 metres and a dis at 45 metres. The cable at 9 metres shows scuffing and possible heat damage to the outside of the cable but no major visible damage - at about 15 metres the cable passes into thick brambles and bushes up to the bridge - the cable enters conduit about a metre from the bridge and has what could be minor rat bite damage as it enters the conduit. 45 metres of cable from DP3X puts the dis at about 3 metres onto the bridge in the conduit. Not good news. Needs more thinking time and hope it doesn’t get worse any time soon.

12 April: Posted by Rick

Rick and Janet at Lydney junction today to test the SPT’s and the PA ready for Saturday (open to the public). All ok.

14 April: Posted by Rick

Rick Janet Peter and Charles in today. Peter initially working at Parkend. Charles and Janet testing Norchard SPTs. SPT2 faulty. Rick checking Norchard clocks. Peter returned to Norchard just as 9681 reported no internet connection. Initially caused by no power - power back on in cafe but still no internet. At the end of the day still no internet to 9681 and SPT 2 still faulty.

16 April: Posted by Rick

Rick Janet and Peter in today. Rick and Janet repaired another break in the dropwire to SPT 2 - now working OK. Peter working on faulty internet at 9681 - assisted by Rick Janet and Bernard Baldwin. After much muttering we found the cable to the socket inside 9681 had been severely chewed by vermin. No time to renew cable so temporary working on wi-fi from Stationmasters office. Another temporary fix added to the list of renewal work.

17 April: Posted by Rick

Rick and Peter in at Parkend on preparation work for the data link for the axle counter. Roger Phelps reported the Lydney junction clock showing incorrect time. Adrian Copley reported the SPT at the south end of the platform faulty.

18 April: Posted by Rick

Rick visited Lydney junction and corrected the clock in the signal box and then rechecked the SPT south end of platform. Tested ok with signalman ( see entry for 12 th April ).

21 April: Posted by Rick

Rick Janet Peter and welcome back Charles today. Charles checking electrical boxes at Norchard for access codes or key availability. Peter working at Parkend on axle counter cct. Rick and Janet changed faulty platform phone at Parkend and then returned to Norchard to check stores in caboose.

23 April: Posted by Rick

Rick attended Norchard at 10.30 after Steve Harris ( S and T ) reported two S and T ccts not working probably due to our cable fault ! No information given on which ccts, but faulty between the abutment and Norchard signal box. The only cable fault outstanding at Norchard is the data link to 9681 and no other ccts involved. Rick checked the records and then opened the cabinet near the abutment and checked the S and T ccts within. Lo and behold there were two broken connections on two separate jumper wires. Wires reterminated and then Steve Harris informed and asked to test if the ccts were working and suggested that S and T tested their ccts better in future.

24 April: Posted by Rick

Peter and Rick at Parkend this morning. While we were sorting out what progress we were making, Dave Collins asked us if we could arrange for an SPT on the UP platform at Parkend. Peter and Rick did a quick survey and measure up and then sent an initial diagram and estimate to Dave Collins. Required for new signal box becoming operational.

28 April: Posted by Rick

Rick Janet Peter and Charles in today. Charles continued checking availability for accessing electrical equipment at Norchard ready for the 2021 electrical site test at Norchard. Peter continued working at Parkend. Rick and Janet went to Parkend and collected some wiring connection blocks suitable to fit at Parkend DP1E where we hope to fit railway slider connections to facilitate ease of fault finding. The platform phone was checked and was working ok.
On our return to Norchard Bob Hawker met us and gave us back our handle for the dropwire dispenser which he had arranged to be repaired (due to a slight accident with a locomotive). A big thank you to Derek Preest for such a good repair.

4 May: Posted by Rick

Rick Janet and Peter in today. All three involved fitting a safety shelf to support the data cabinet in the cafe (appears to be falling off the wall as the glue is failing). The large clock in the main shop area with the two clock faces were showing different times and were were realigned. Alex Davies was on site and agreed to arrange transportation of the 20 pr cable to Parkend on Thursday ready for laying out between the ground frame and Parkend exchange.
Green ducting was recovered from under the upper platform and laid out ready for digging in during the work to extend the rail line in front of the telecoms storage container. For future cable provision around Norchard.

5 May: Posted by Rick

Rick Charles and Peter in today. Lloyd Townsend arrived and agreed to start electrical testing on 17th May. Charles still checking access to electrical equipment. Peter at Parkend. Rick checking cable and cable jacks ready for move to Parkend.
Peter reported that the power to Parkend exchange had been turned off and the batteries were virtually flat but power was back on and the voltage was coming back and by lunchtime it appeared to have stabilised. Let’s hope there is no permanent damage to the batteries.

6 May: Posted by Rick

Cable and jacks delivered to Parkend and covered by tarpaulins.

12 May: Posted by Rick

Rick Janet and Peter in today at Parkend. Although we had hoped to lay the 20 pr cable on Saturday the weather was so bad we gave it a miss. Today we were successful and the 20 pr cable was laid in the troughing from new Parkend DP 1 (ground frame )to Parkend exchange. Due to the delay from Saturday we were a bit short handed and had to rely on press ganging whoever was around to assist us but fortunately most were willing to help if only for a short period.

23 May: Posted by Rick

Rick received an e mail from Alex Davies who had received an e mail about cables tied to the fence at Lydney Fire Station. As the fence was due for replacement the council wanted us to take them off the fence. Rick visited Lydney Town Station and verified that the cables in question belonged to the DFR. An urgent job for Wednesday!


Return to the Diary Menu


mainmenu

Page provided by John Bathgate

This page was last updated on
18th June 2021